The sum of two fractions is 2 over 3. One fraction is 1 over 2. What is the other fraction

Answer by Alan3354(69443)

2 over 3? that's 2/3
-----
2/3 = 1/2 + the other fraction
the other fraction = 2/3 - 1/2